# On the Impact of Uncertainty and Calibration on Likelihood-Ratio Membership Inference Attacks

### v2

@@ Abstract
In a membership inference attack (MIA), an attacker exploits the overconfidence exhibited by typical machine learning models to determine whether a specific data point was used to train a target model. In this paper, we analyze the performance of the state-of-the-art likelihood ratio attack (LiRA) within an information-theoretical framework that allows the investigation of the impact of the aleatoric uncertainty in the true data generation process, of the epistemic uncertainty caused by a limited training data set, and of the calibration level of the target model. We compare three different settings, in which the attacker receives decreasingly informative feedback from the target model: confidence vector (CV) disclosure, in which the output probability vector is released; true label confidence (TLC) disclosure, in which only the probability assigned to the true label is made available by the model; and decision set (DS) disclosure, in which an adaptive prediction set is produced as in conformal prediction. We derive bounds on the advantage of an MIA adversary with the aim of offering insights into the impact of uncertainty and calibration on the effectiveness of MIAs. Simulation results demonstrate that the derived analytical bounds predict well the effectiveness of MIAs.
@@ I-A Context and Motivation
Furthermore, for a black-box attacker, epistemic uncertainty creates the additional complication of increasing the uncertainty about the actual model parameters of the target model. In fact, the variability in the model parameter space associated with an unknown training data set increases with smaller data set sizes, since larger data sets tend to all yield models that minimize the underlying population loss.
@@ I-B State of the Art
Reference [2] introduced the likelihood ratio attack (LiRA), in which Gaussian distributions are fit to the TLC generated by shadow models trained to imitate models trained with and without the candidate data point. LiRA is then conducted through a parametric likelihood ratio test. Reference [15] improved LiRA by adding adversarial noise to the input to amplify the output gap between the two classes of shadow models (see also [16]).
@@
Theoretical analysis of membership inference attacks: Numerous studies have explored the link between model calibration and privacy risks via experiments [1, 2, 5] or via scenario-specific theoretical analyses [3, 4, 12, 22, 23]. Notably, a line of work is motivated by the opposite nature of the goal of MIAs – revealing a training data point – and of differential privacy (DP) mechanisms – obscuring the presence or absence of a training data point [24]. Based on this observation, upper bounds on the attacker’s advantage [3, 4] and positive predictive value [12] were studied for DP-protected models. These works focus on the theoretical relationship between the DP metrics and MIA success measures. Furthermore, reference [22] established universal bounds on the success rate of MIAs as a function of the target model’s generalization gap, while assuming a white-box attacker. Finally, results that target symmetric and redundancy-invariant algorithms were presented in [23].
@@ I-C Contributions and Organization
In this paper, we introduce a unified theoretical framework to analyze the performance of MIA under CV, TLC, and DS disclosures, while assuming black-box attacks as in [2, 1] and focusing on the impact and interplay of calibration and predictive uncertainty. The main contributions of this paper are summarized as follows.
  • As illustrated in Fig. 1, treating MIA design for black-box models as a hypothesis-testing problem, we introduce a general information-theoretical framework to analyze the attacker’s advantage for CV, TLC, and DS disclosure settings. Specifically, as seen in Fig. 2, the proposed methodology addresses LiRA-style attacks built from shadow models that imitate the behavior of the target model to inputs inside and outside the training set.
@@
The remainder of this paper is organized as follows. In Sec. II, we outline the problem, review the LiRA framework, and introduce relevant performance metrics. Sec. III reviews a performance bound for MIA, and defines aleatoric uncertainty and calibration. In Sec. IV, we analyze the performance of MIA given CV, TLC, and DS observations. Sec. V empirically evaluates the performance of MIA, validating our theoretical findings. Sec. VI concludes the paper.
@@ II Problem Definition
[…]
II-A Membership Inference Attack as Hypothesis Testing
Focusing on a $K$-class classification problem, the target model $p(y|x,\theta)$ is a conditional probability distribution over the label variable $y\in\left\{0,1,\ldots,K-1\right\}$ given input $x$ that is parameterized by a vector $\theta$. The model parameter vector $\theta$ is optimized by the target model on the basis of a training algorithm
$\theta_{\mathcal{D}^{\mathrm{tr}}}=\mathcal{T}(\mathcal{D}^{\mathrm{tr}}),$ (1)
where function $\mathcal{T}(\cdot)$ may incorporate any randomness arising from stochastic steps such as mini-batch selection. The training data set $\mathcal{D}^{\mathrm{tr}}$ consists of pairs $(x,y)$ generated from the ground-truth distribution $p^{*}(x,y)$.
As mentioned in [2], we assume that the attacker knows the data generation distribution $p^{*}(x,y)$, as well as the training algorithm $\mathcal{T}(\cdot)$ in (1). However, it does not have access to the training data set $\mathcal{D}^{\mathrm{tr}}$ and to the trained parameter vector $\theta_{\mathcal{D}^{\mathrm{tr}}}$. Furthermore, the attacker can query once the target model with input $x_{0}$ of interest.
The information released by the target model is a function of the confidence probability vector $p(y|x_{0},\theta_{\mathcal{D}^{\mathrm{tr}}})$ output by the model in response to a query with input $x_{0}$. We write the confidence probability vector as
@@ II-A Membership Inference Attack as Hypothesis Testing
Based on the available information on distribution $p^{*}(x,y)$, training algorithm $\mathcal{T}(\cdot)$, and output $O(p)$ of the model (3), (4), or (5), in LiRA, the attacker aims at distinguishing two hypotheses
@@ II-B Likelihood Ratio Attack
In order to construct the test variable $T(O)$, in this paper, we focus on the state-of-the-art likelihood ratio attack (LiRA). As illustrated in Fig. 2, LiRA estimates the distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ of the observation $O$ during a preliminary phase. It is emphasized that the original LiRA focused solely on TLC outputs (4), and assumed Gaussian distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ to obtain a practical algorithm. In this paper, we study a more general framework, allowing for any of the outputs $O$ in (3)–(5), with the aim of deriving analytical insights.
@@
Distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ represent the probability density functions of the observations $O$ under the hypotheses $\mathcal{H}^{\mathrm{out}}$ and $\mathcal{H}^{\mathrm{in}}$, respectively. Accordingly, the randomness captured by the distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ reflects the inherent uncertainty of the attacker about the training data set $\mathcal{D}^{\mathrm{tr}}$ used to optimize the target model as per (1). For the analysis in what follows, we assume $L$ to be large enough that distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ are correctly estimated by the attacker. This assumption represents a meaningful worst-case condition for the target model, since $L$ is only limited by the computational complexity of the attacker.
@@ III-C Calibration
While the true probability of the class $y_{0}$ is $p_{0}^{*}$, the model assigns to it a probability $p_{0}$. A well-calibrated model is expected to output a confidence probability $p_{0}$ that is close to the ground-truth probability $p_{0}^{*}$ [27, 21]. However, machine learning models tend to be overconfident, assigning a probability $p_{0}$ larger than $p_{0}^{*}$, especially when trained with the example $(x_{0},y_{0})$ being evaluated. In fact, it is well known that the power of MIA depends on the level of overconfidence of the target model [2, 16, 1, 5, 10]. A more overconfident model tends to assign a larger probability $p_{0}$, making it easier to detect the presence of example $(x_{0},y_{0})$ in the training data set.
@@ IV-A Modeling the Output Distributions
with $z\in\left\{\mathrm{out},\mathrm{in}\right\}$. The variance (IV-A) accounts for the dependence of the trained model as a function of the training sets drawn from the ground-truth distribution $p^{*}(x,y)$. This model-level uncertainty, known as epistemic uncertainty [6], decreases as the size $N^{\mathrm{tr}}$ of the training data set increases. In fact, in the regime $N^{\mathrm{tr}}\rightarrow\infty$, the training data sets are fully representative of the distribution $p^{*}(x,y)$, and the probabilities $p_{k}$ tend to the ground-truth values $p^{*}_{k}$ [29, 30]. Here, since we target black-box attacks, for which model parameters are unavailable, we measure the epistemic uncertainty indirectly through the variance observed in the model output distribution.
The possibility to decrease the epistemic uncertainty is in stark contrast to the aleatoric uncertainty introduced in the previous section, which is due to the inherent randomness of the data, and thus it does not decrease as the training set size increases. It is therefore possible to simultaneously have a high epistemic uncertainty due to limited data and a low aleatoric uncertainty for “easy” examples; and, vice versa, a high aleatoric uncertainty for “hard” examples and a low epistemic uncertainty due to access to a large training set.
@@
Furthermore, by (28), since the data set size $N^{\mathrm{tr}}$ is the same for both data sets $\mathcal{D}^{\mathrm{out}}$ and $\mathcal{D}^{\mathrm{in}}$, we assume the sum in (28) to be the same for both distributions $f^{\mathrm{out}}(p)$ and $f^{\mathrm{in}}(p)$.
@@
Note that, in order to ensure the inequalities $\gamma^{z}_{k}>0$ with $z\in\left\{\mathrm{out},\mathrm{in}\right\}$ and $k=0,1,\ldots,K-1$, we have the inequality $\Delta/(1+\Delta)<\epsilon_{a}<1$. Equations (29)–(30) indicate that the average confidence generated by the model in the true label $y_{0}$ decreases with the aleatoric uncertainty $\epsilon_{a}$ and with epistemic uncertainty $\epsilon_{e}$, while also increasing with the relative calibration error $\Delta$ for the model trained with $(x_{0},y_{0})$.
@@ IV-B Confidence Vector Disclosure
Proof: The proof of (1) is detailed in Appendix -A.
@@ IV-D Decision Set Disclosure
Proof: The proof of (38) is detailed in Appendix -B.
@@ V-A On the Dirichlet Assumption
Fig. 5 illustrates the probability density functions of the confidence level $p_{0}$ for the target sample $(x_{0},y_{0})$, along with the corresponding empirical distribution obtained from the fitted Dirichlet distributions for the models trained on $\mathcal{D}^{\mathrm{out}}$ (left) and $\mathcal{D}^{\mathrm{in}}$ (right). The model is observed to be overconfident when trained by including the sample $(x_{0},y_{0})$. The empirical distributions closely align with the fitted Dirichlet distributions, confirming the flexibility of the working assumption of Dirichlet distributions for the confidence vectors. Throughout the rest of this section, we adopt this assumption, which is further validated in Appendix -C.
@@ VI Conclusions
This paper has introduced a theoretical framework to analyze the attacker’s advantage in MIAs as a function of the model’s calibration, as well as of the aleatoric and epistemic uncertainties. Following the state-of-the-art LiRA-style attacks, we regarded MIA as a hypothesis-testing problem; and we considered three typical types of observations that can be acquired by the attacker when querying the target model: confidence vector (CV), true label confidence (TLC), and decision set (DS). By making the flexible modeling assumption that the output confidence vectors follow the Dirichlet distributions, thus accounting for both aleatoric and epistemic uncertainties, we derived upper bounds on the attacker’s advantage. The bounds yield analytical insights on the impact of the relative calibration error, aleatoric uncertainty, and epistemic uncertainty on the success of MIAs. For DS disclosure, we studied for the first time the impact of the threshold and of randomization on the effectiveness of MIA.
@@
-A Proof of Proposition 1
@@
-B Proof of Proposition 3
We leverage the strong data processing inequality [38, Proposition II.4.10][39] to obtain the upper bound of the MIA advantage given DS observations as
@@
where we have used Lemma 2 and (-B).
-C Additional Validation of the Dirichlet Assumption
@@
-D Performance in High TNR Regime

### v3

@@ Abstract
In a membership inference attack (MIA), an attacker exploits the overconfidence exhibited by typical machine learning models to determine whether a specific data point was used to train a target model. In this paper, we analyze the performance of the likelihood ratio attack (LiRA) within an information-theoretical framework that allows the investigation of the impact of the aleatoric uncertainty in the true data generation process, of the epistemic uncertainty caused by a limited training data set, and of the calibration level of the target model. We compare three different settings, in which the attacker receives decreasingly informative feedback from the target model: confidence vector (CV) disclosure, in which the output probability vector is released; true label confidence (TLC) disclosure, in which only the probability assigned to the true label is made available by the model; and decision set (DS) disclosure, in which an adaptive prediction set is produced as in conformal prediction. We derive bounds on the advantage of an MIA adversary with the aim of offering insights into the impact of uncertainty and calibration on the effectiveness of MIAs. Simulation results demonstrate that the derived analytical bounds predict well the effectiveness of MIAs.
@@ I-A Context and Motivation
Furthermore, an attacker without access to the target model weights is characterized by epistemic uncertainty, i.e., by uncertainty about the actual model parameters of the target model [10]. In fact, the variability in the model parameter space associated with an unknown training data set increases with smaller data set sizes, since larger data sets tend to all yield models that minimize the underlying population loss.
@@ I-B State of the Art
Reference [2] introduced a type of likelihood ratio attack (LiRA), in which Gaussian distributions are fitted to the TLC generated by shadow models trained to imitate models trained with and without the candidate data point. LiRA is then conducted through a parametric likelihood ratio test. Reference [16] improved LiRA by adding adversarial noise to the input in order to amplify the output gap between the two classes of shadow models. Further enhancements include setting sample-specific thresholds [17] and the use of pairwise likelihood ratio tests calibrated through a Bayesian method [18].
@@
[…]
Theoretical analysis of membership inference attacks: Previous work has explored the impact of various factors on privacy risks via experiments [1, 2, 5, 24, 25] or via scenario-specific theoretical analyses [3, 4, 13, 26, 27, 28, 29]. While some studies have empirically identified factors such as data complexity, intra-class variance, number of classes, data entropy, model type, and model fairness as influential to MIA performance [24, 25], these factors can ultimately be viewed as indirect manifestations of uncertainty and calibration. Given the multitude of such factors, our work aims to rigorously explore the fundamental impact of uncertainty and calibration on MIA performance through theoretical analysis.
Notably, a line of work is motivated by the opposite nature of the goal of MIAs – revealing a training data point – and of differential privacy (DP) mechanisms – obscuring the presence or absence of a training data point [30]. Based on this observation, upper bounds on the attacker’s advantage [3, 4] and positive predictive value [13] were studied for DP-protected models. These works focus on the theoretical relationship between the DP metrics and MIA success measures. Furthermore, reference [26] established universal bounds on the success rate of MIAs as a function of the target model’s generalization gap, while assuming a white-box attacker. Results that target symmetric and redundancy-invariant algorithms were also discussed in [27]. Further theoretical analyses have explored the vulnerability to MIA across population subgroups [28], as well as the impact of generalization [29].
@@ I-C Contributions and Organization
In this paper, we introduce a unified theoretical framework to analyze the performance of MIA under CV, TLC, and DS disclosure modes, assuming that the attacker follows the membership inference game studied in [3, 13, 2, 17, 18] (see Definition 1 in Sec. II-A). Rather than proposing a new attack, our objective is to provide a theoretical analysis of existing LiRA-style attacks, focusing on the impact and interplay of calibration and predictive uncertainty. The main contributions of this paper are summarized as follows.
  • As illustrated in Fig. 1, treating MIA design as a hypothesis-testing problem, we introduce a general information-theoretical framework to analyze the attacker’s advantage for CV, TLC, and DS disclosure settings. Specifically, as seen in Fig. 2, our methodology analyzes LiRA-style attacks [2] for a wide range of disclosure settings, beyond the case of TLC disclosure considered in prior work. LiRA-style attacks use shadow models to imitate the behavior of the target model for inputs inside and outside the training set.
@@
The remainder of this paper is organized as follows. In Sec. II, we outline the problem, review the LiRA framework, and introduce relevant performance metrics. Sec. III reviews a performance bound for MIA, and defines aleatoric uncertainty and calibration. In Sec. IV, we analyze the performance of MIA given CV, TLC, and DS observations. Sec. V presents simulation-based results to evaluate the performance of MIA, validating our theoretical findings. Sec. VI concludes the paper.
@@ II Problem Definition
II-A Setting
We focus on a $K$-class classification problem in which the target model belongs to a parameterized model class
=$\mathcal{F}=\left\{p(y|x,\theta):\theta\in\mathbb{R}^{d}\right\},$ (1)
consisting of conditional probability distributions over the label variable $y\in\left\{0,1,\ldots,K-1\right\}$ given input $x$ that is parameterized by a vector $\theta$. Let $\mathcal{T}(\cdot)$ denote a training algorithm that may incorporate any randomness arising from stochastic steps such as mini-batch selection, and $p^{*}(x,y)$ be the ground-truth data distribution. As in prior art [3, 13, 2, 17, 18], we define the membership inference game as follows.
Definition 1 (Membership Inference Game).
A membership inference game consists of the following steps:
  • 1)
    A training data set $\mathcal{D}^{\mathrm{tr}}$ is generated by sampling data points in an independent and identically distributed (i.i.d.) way from the ground-truth distribution $p^{*}(x,y)$. This training data set is used to train the target model $p(y|x,\theta_{\mathcal{D}^{\mathrm{tr}}})$, with trained parameters obtained from the training algorithm as $\theta_{\mathcal{D}^{\mathrm{tr}}}=\mathcal{T}(\mathcal{D}^{\mathrm{tr}})$.
  • 2)
    A bit $b\in\left\{0,1\right\}$ is drawn uniformly at random. If $b=1$, a data point $(x_{0},y_{0})$ is selected from the training set $\mathcal{D}^{\mathrm{tr}}$, while, if $b=0$, a new data point $(x_{0},y_{0})\sim p^{*}(x,y)$ is generated from the ground-truth distribution in a manner independent of the training set.
  • 3)
    Given the pair $(x_{0},y_{0})$, the attacker estimates the binary variable $b$ by leveraging its knowledge of the ground-truth distribution $p^{*}(x,y)$, of the model class $\mathcal{F}$, and of the training algorithm $\mathcal{T}(\cdot)$.
This game makes the worst-case assumption that the adversary knows the data generation distribution $p^{*}(x,y)$, so that it can generate data from it. In practice, this situation is approximately valid as many attacks require access to samples from the ground-truth distribution [1, 3, 13, 11, 2, 17, 18, 24]. From these samples, the data generation distribution can be estimated using techniques such as model-based synthesis and statistics-based synthesis [1]. The game also posits that the attacker can use the same model architecture $\mathcal{F}$ and training algorithm $\mathcal{T}(\cdot)$ as the target model. This may be the case if the adversary can leverage the same machine learning-as-a-service provider that builds the target model [1, 11], or if it performs model extraction to approximate the target model’s behavior [31]. Importantly, the attacker does not have access to the target model’s training data set $\mathcal{D}^{\mathrm{tr}}$ or to its trained parameter vector $\theta_{\mathcal{D}^{\mathrm{tr}}}$.
II-B Likelihood Ratio Attack
In this paper we study a membership inference attack (MIA) that follows the assumptions of the membership inference game in Definition 1. The attacker observes the output of the target model for the given input $x_{0}$. The information released by the target model is a function of the confidence probability vector $p(y|x_{0},\theta_{\mathcal{D}^{\mathrm{tr}}})$ output by the model in response to a query with input $x_{0}$. We write the confidence probability vector as
@@ II-B Likelihood Ratio Attack
with $p_{k}\in[0,1]$ and $\sum_{k=0}^{K-1}p_{k}=1$. Without loss of generality, we reorder the entries so that the first element, $p_{0}$, corresponds to the probability assigned by the model to the true label $y_{0}$.
@@
Based on the output $O(p)$ of the model (3), (4), or (5), in LiRA, the attacker aims at distinguishing two hypotheses
@@
In order to construct the test variable $T(O)$, in this paper, we focus on the likelihood ratio attack (LiRA) [2], which follows the membership inference game in Definition 1. We refer to Sec. VI for some discussions on LiRA-style attacks with less powerful adversaries [16, 17, 18]. As illustrated in Fig. 2, LiRA estimates the distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ of the observation $O$ during a preliminary phase. It is emphasized that the original LiRA focused solely on TLC outputs (4), and assumed Gaussian distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ to obtain a practical algorithm. In this paper, we study a more general framework, allowing for any of the outputs $O$ in (3)–(5), with the aim of deriving analytical insights.
@@
Distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ represent the probability density functions of the observations $O$ under the hypotheses $\mathcal{H}^{\mathrm{out}}$ and $\mathcal{H}^{\mathrm{in}}$, respectively. Accordingly, the randomness captured by the distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ reflects the inherent uncertainty of the attacker about the training data set $\mathcal{D}^{\mathrm{tr}}$ used to optimize the target model using algorithm $\mathcal{T}(\cdot)$. For the analysis in what follows, we assume $L$ to be large enough that distributions $f^{\mathrm{out}}(O)$ and $f^{\mathrm{in}}(O)$ are correctly estimated by the attacker. This assumption represents a meaningful worst-case condition for the target model, since $L$ is only limited by the computational complexity of the attacker.
@@ III-C Calibration
While the true probability of the class $y_{0}$ is $p_{0}^{*}$, the model assigns to it a probability $p_{0}$. A well-calibrated model is expected to output a confidence probability $p_{0}$ that is close to the ground-truth probability $p_{0}^{*}$ [34, 23]. However, machine learning models tend to be overconfident, assigning a probability $p_{0}$ larger than $p_{0}^{*}$, especially when trained with the example $(x_{0},y_{0})$ being evaluated. In fact, it is well known that the level of overconfidence of the target model is a significant factor influencing the power of MIA [2, 17, 1, 5, 11]. A more overconfident model tends to assign a larger probability $p_{0}$, making it easier to detect the presence of example $(x_{0},y_{0})$ in the training data set.
@@ IV-A Modeling the Output Distributions
with $z\in\left\{\mathrm{out},\mathrm{in}\right\}$. The variance (IV-A) accounts for the dependence of the trained model as a function of the training sets drawn from the ground-truth distribution $p^{*}(x,y)$. This model-level uncertainty, known as epistemic uncertainty [6], decreases as the size $N^{\mathrm{tr}}$ of the training data set increases and as the model’s capacity to learn the underlying data distribution improves. In fact, in the regime $N^{\mathrm{tr}}\rightarrow\infty$ where the training data sets are fully representative of the distribution $p^{*}(x,y)$, and assuming the model has sufficient capacity, the learned probabilities $p_{k}$ tend to the ground-truth values $p^{*}_{k}$ [37, 38]. With a finite training set, an attacker following Definition 1 is characterized by uncertainty on the model, which we account for in our analysis through the variance of the model output distribution [39].
The possibility to decrease the epistemic uncertainty is in stark contrast to the aleatoric uncertainty introduced in the previous section, which is due to the inherent randomness of the data, and thus it does not decrease as the training set size or the model capacity increases. It is therefore possible to simultaneously have a high epistemic uncertainty due to limited data and a low aleatoric uncertainty for “easy” examples; and, vice versa, a high aleatoric uncertainty for “hard” examples and a low epistemic uncertainty due to access to a large training set.
@@
Furthermore, by (28), since both the data set size $N^{\mathrm{tr}}$ and the model class $\mathcal{F}$ in (1) are the same for both data sets $\mathcal{D}^{\mathrm{out}}$ and $\mathcal{D}^{\mathrm{in}}$, we assume the sum in (28) to be the same for both distributions $f^{\mathrm{out}}(p)$ and $f^{\mathrm{in}}(p)$.
@@
Please refer to Appendix -A for a derivation of (29) and (30). Note that, in order to ensure the inequalities $\gamma^{z}_{k}>0$ with $z\in\left\{\mathrm{out},\mathrm{in}\right\}$ and $k=0,1,\ldots,K-1$, we have the inequality $\Delta/(1+\Delta)<\epsilon_{a}<1$. Equations (29)–(30) indicate that the average confidence generated by the model in the true label $y_{0}$ decreases with the aleatoric uncertainty $\epsilon_{a}$ and with epistemic uncertainty $\epsilon_{e}$, while also increasing with the relative calibration error $\Delta$ for the model trained with $(x_{0},y_{0})$.
@@ IV-B Confidence Vector Disclosure
Proof: The proof of (1) is detailed in Appendix -B.
@@ IV-D Decision Set Disclosure
Proof: The proof of (38) is detailed in Appendix -C.
@@ V-A On the Dirichlet Assumption
Fig. 5 illustrates the probability density functions of the confidence level $p_{0}$ for the target sample $(x_{0},y_{0})$, along with the corresponding empirical distribution obtained from the fitted Dirichlet distributions for the models trained on $\mathcal{D}^{\mathrm{out}}$ (left) and $\mathcal{D}^{\mathrm{in}}$ (right). The model is observed to be overconfident when trained by including the sample $(x_{0},y_{0})$. The empirical distributions closely align with the fitted Dirichlet distributions, confirming the flexibility of the working assumption of Dirichlet distributions for the confidence vectors. Throughout the rest of this section, we adopt this assumption, which is further validated in Appendix -D.
@@ VI Conclusions
This paper has introduced a theoretical framework to analyze the attacker’s advantage in MIAs as a function of the model’s calibration, as well as of the aleatoric and epistemic uncertainties. Following the theoretical framework of LiRA-style attacks, we regarded MIA as a hypothesis-testing problem; and we considered three typical types of observations that can be acquired by the attacker when querying the target model: confidence vector (CV), true label confidence (TLC), and decision set (DS). By making the flexible modeling assumption that the output confidence vectors follow the Dirichlet distributions, thus accounting for both aleatoric and epistemic uncertainties, we derived upper bounds on the attacker’s advantage. The bounds yield analytical insights on the impact of the relative calibration error, aleatoric uncertainty, and epistemic uncertainty on the success of MIAs. For DS disclosure, we studied for the first time the impact of the threshold and of randomization on the effectiveness of MIA.
Our theoretical analysis is formulated based on the LiRA-style attacks that assume the worst-case scenario defined by the membership inference game in Definition 1. Accordingly, the attacker has accurate knowledge of the data distribution, as well as of the target model’s architecture and training algorithm. We also do not impose constraints on computational resources, enabling the attacker to train enough shadow models to accurately estimate the output distributions for models trained with and without the target data point. This setting represents the most favorable conditions for the attacker, providing an upper bound on the MIA advantage.
Given that the derived bounds assume a worst-case adversary, the analysis in this paper is also applicable to recent LiRA-style atacks that inherently limit attack performance. These include settings with data shifts unknown to the attacker, as well as computational constraints [16, 17, 18]. Notably, reference [18] uses a Bayesian method to provide a lightweight version of LiRA-style attack, requiring only one or two shadow models.
For practical guidance on the design of protection strategies against LiRA-based MIAs derived from our theoretical analysis, please see Appendix -F.
@@
-A Derivation of the Dirichlet Parameters in (29) and (30)
Equations (29) and (30), which express the Dirichlet parameters for the distributions of $f^{\mathrm{out}}(p)$ and $f^{\mathrm{in}}(p)$ as functions of the aleatoric uncertainty $\epsilon_{a}$ in (23), epistemic uncertainty $\epsilon_{e}$ in (28), and relative calibration error $\Delta$ in (24), are obtained as follows.
By substituting the aleatoric uncertainty $\epsilon_{a}$ from (23) into the relative calibration error $\Delta$ in (24), the expected confidence $\mathbb{E}_{f^{\mathrm{in}}}\left[p_{0}\right]$ can be expressed in terms of parameters $\epsilon_{a}$ and $\Delta$. Then, substituting this expected confidence and the epistemic uncertainty $\epsilon_{e}$ from (28) into the average confidence level (26) with $z=\mathrm{in}$ and $k=0$, the parameter $\gamma^{\mathrm{in}}_{0}$ can be represented as a function of parameters $\Delta$, $\epsilon_{a}$, and $\epsilon_{e}$. Using the assumption $\gamma^{\mathrm{in}}_{1}=\ldots=\gamma^{\mathrm{in}}_{K-1}$ for all other candidate classes, we get $\gamma^{\mathrm{in}}_{k}$ for $k=1,2\ldots,K-1$ using (28). For the distribution $f^{\mathrm{out}}(p)$, the parameters $\gamma^{\mathrm{out}}_{k}$ can be obtained by setting $\Delta=0$ in the expression for $\gamma^{\mathrm{in}}$ in (30).
-B Proof of Proposition 1
@@
-C Proof of Proposition 3
We leverage the strong data processing inequality [47, Proposition II.4.10][48] to obtain the upper bound of the MIA advantage given DS observations as
@@
where we have used Lemma 2 and (-C).
-D Additional Validation of the Dirichlet Assumption
@@
-E Performance in High TNR Regime